6x-9y=25.5
7y-3x=-14
solve by elimination, to include solution, if any & ordered pair.

Answer by ewatrrr(24785)

 
Hi
6x-9y=25.5
7y-3x=-14
6x-9y = 25.5
-3x+7y = -14
6x-9y = 25.5
-6x+14y = -28
5y = -2.5
y = -.5 and x = 3.5 (-3.5 + 14)/3
CHECKING our Answer***
6*3.5 -9*-.5 = 21 +4.5 = 25.5